OpenStudy (anonymous):

What are the diameter and radius of a circle with a circumference of 28π m? A. d = 28 m r = 14 m B. d = 14 m r = 28 m C. d = 14 m r = 7 m D. d = 7 m r = 14 m

jimthompson5910 (jim_thompson5910):

The circumference of a circle is C = 2*pi*r In this case, C = 28pi (given)

jimthompson5910 (jim_thompson5910):

plug that in to get C = 2*pi*r 28pi = 2*pi*r now solve for r and tell me what you get

OpenStudy (anonymous):

I think u would dive 28 into3 but then that does not work because u would then add up to 9.33333333

jimthompson5910 (jim_thompson5910):

Here's how you solve for the radius r 28pi = 2*pi*r 28pi/pi = 2*pi*r/pi ... Divide both sides by pi. 28 = 2*r 28/2 = 2r/2 ... Divide both sides by 2. 14 = r r = 14

jimthompson5910 (jim_thompson5910):

To find the diameter d, you use the formula d = 2*r basically the diameter is twice as long as the radius